OpenStudy (jbrocato22):

The surface area of a right circular cylinder of height 4 feet and radius r feet is given by S(r)=2πrh+2πr2. Find the instantaneous rate of change of the surface area with respect to the radius, r, when r = 4.

OpenStudy (amity):

ds/dr=2pie h+4pie r=pie(2h+4r)=3.14(2•4+4•4)=3.14•24=75.36feet^2/feet

OpenStudy (amity):

ds/dr=2pie h+4pie r=pie(2h+4r)=3.14(2•4+4•4)=3.14•24=75.36feet^2/feet